OpenStudy (anonymous):

Y=mx + b 6= 6/13(-3) +b Solve for B ???

OpenStudy (anonymous):

B= 7.4

OpenStudy (anonymous):

0r 96/13 to be precise

OpenStudy (anonymous):

Can you explain to me how you worked this problem out?

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

I have to find y=mx +b my slope is 6/13... (-3,6) when I plug it in this is what I get 6= 6/13(-3) + b now I have to solve for b???

OpenStudy (anonymous):

Im confused because of the fraction

OpenStudy (anonymous):

y=mx+b <------slope intercept formula 6=6/13( -3) + b <----your equation 6= -18/13 + b <---- you have to multiply 6/13 times -3 which gives you -18/13 now you have to add -18/3 to 6 which gives you 60/13 I mad a mistake sorry....the answer is 4.61538 which rounds to 4.6

OpenStudy (anonymous):

OK. you added -18/3 i thought its suppose to be -18/13.. ??

OpenStudy (anonymous):

How did you get 60/13??